HETAV PATEL

Grade 11,

a submarine goes 30 km upstream and 44 km downstream in 10 hrs it can also go 40 km up stream and 55 km down stream in 13 hrs find the speed of the boat in still water and speed of stream

a submarine goes 30 km upstream and 44 km downstream in 10 hrs it can also go 40 km up stream and 55 km down stream in 13 hrs find the speed of the boat in still water and speed of stream

Grade:9

1 Answers

Sakshi
7 years ago
Speed of boat in still water is 8and speed of stream is 3 km/hr let us suppose that speed of boat instill water is x and speed of stream i y equations are. 30/x-y +44/x+y=10...(1). And 40/x-y+55/x+y=13...(2)